# Hesitation and Tolerance in Recommender Systems

### v1

@@ 7. Conclusion
Our user study revealed that hesitation frequently occurs when users deliberate over whether to engage with recommended items. This hesitation often signals ambivalence, as items that provoke it are more likely to align with a decline in overall user interest. Interestingly, hesitation is often accompanied by deeper engagement, as users may explore the item further in an effort to resolve their uncertainty. However, when the item ultimately fails to meet their expectations, this extended engagement transforms into what we define as tolerance behavior. Tolerance behavior is characterized by an initial sequence of positive actions, such as clicks or extended viewing, which eventually shift to negative actions like disengagement or dissatisfaction. This progression makes tolerance behavior a clear and actionable negative signal for optimizing recommender systems.
By explicitly incorporating tolerance behavior into our optimization framework, we tested its impact on a large-scale industrial video platform. Our A/B testing demonstrated that addressing tolerance behavior led to a significant boost in user retention rates. This improvement was achieved with minimal implementation costs, i.e., simply changing the sample labels during model optimization, highlighting the practical value and scalability of this approach. Furthermore, this method provides a deeper understanding of user behavior dynamics, emphasizing the importance of moving beyond traditional engagement metrics to consider more nuanced patterns of interaction. These insights can help drive the development of smarter, more user-centric recommendation strategies that foster long-term satisfaction and loyalty.
We acknowledge that our survey design is limited two recommendation scenarios, e-commerce and content browsing or more specifically short-video sharing. Our online experiments are limited to the latter scenario only. Whether the findings generalize to more recommendation settings remains unclear and more research effert is needed.

### v2

@@ 7. Conclusion
Our study reveals that hesitation is a common part of decision-making in recommender systems. While hesitation signals uncertainty, it often leads users to invest additional effort exploring items — only to end in disappointment. We define this progression as tolerance behavior: interactions that begin with curiosity and engagement but turn into frustration and withdrawal. Tolerance, therefore, offers a powerful signal that current engagement metrics overlook.
By explicitly modeling tolerance in a large-scale video platform, we demonstrated that reframing shallow or regretful interactions as distinct feedback improved next-day retention with minimal engineering cost. This shows that even small adjustments to how we treat user signals can meaningfully shift outcomes, offering both practical scalability and theoretical insight.
More broadly, our findings underscore the need to move beyond clicks and watch time as proxies for “success.” Designing recommender systems that minimize wasted effort and respect user time fosters deeper trust, more sustainable engagement, and ultimately more value for both users and platforms.
It is important to acknowledge that our work is subject to certain limitations. We focused on two domains, e-commerce and short video, and validated our modeling approach in only the latter. Whether tolerance manifests similarly in other contexts, such as news, live streaming, or productivity tools, remains an open question. Future work should also explore longer-term outcomes and richer qualitative measures to capture the lived experience of hesitation and tolerance.